# Density-Matrix Mean-Field Theory

### v1

@@ 3.1 Affleck-Kennedy-Lieb-Tasaki Model
The AKLT model is an exactly solvable system, showcasing non-trivial topological order. Notably, it features spin-$1/2$ (fractional to spin-$1$) edge states and exhibits $4$-fold degeneracy of the ground states for an open chain. This makes the AKLT model a valuable benchmark example.
@@
We apply the DMMFT to the AKLT model, focusing the two-site cluster $c$ and solving the reduced DM $\rho_{c}^{(2)}$ self-consistently according to the mean-field equations as described in Sec. 2.2. Specifically, the reduced DM is calculated over an extended cluster encompassing the left and right nearest neighboring two-site clusters to $c$. The associated Hilbert space is denoted as ~$\tilde{\mathscr{H}}_{\bar{c}}=\tilde{\mathscr{H}}_{l}\otimes\mathscr{H}_{c}\otimes\tilde{\mathscr{H}}_{r}$, where ~$\tilde{\mathscr{H}}_{l,r}\sim\Pi_{c}(\mathscr{H}_{c})$ are the Hilbert subspaces selected by the reduced DM, as per Eq.(13). We choose the cut-off parameter $n_{c}=4$.
In Fig. 1, the spectra of the two-site and single-site reduced DMs are presented. Blue circles represent exact values, while red dots depict DMMFT results. The spectrum of the two-site reduced DM obtained with the DMMFT reasonably aligns with the exact values [Fig. 1(a)], notably capturing the correct degeneracies. The entanglement entropy of the two-site cluster evaluated byDMMFT is $SE_{c,\text{MF}}^{(2)}=1.58$. Although it is slightly larger than the exact value $SE_{c,\text{exact}}^{(2)}=1.3689$, this discrepancy is expected due to the cut-off parameter $n_{c}$ chosen being marginally larger than $\exp(SE_{c,\text{exact}}^{(2)})=3.9310$. Fig. 1(b) displays the spectrum of the single-site reduced DM, where the DMMFT results align excellently with the exact values. The single-site entanglement entropy $SE_{c}^{(1)}=1.0986=\ln(3)$ also matches. Furthermore, the expectation values of two spins within the cluster $\langle S_{1,2}^{\alpha}\rangle=\text{Tr}_{c}(S_{1,2}^{\alpha}\rho_{c})$ both vanish. This implies that the ground states found by DMMFT are not biased towards the semiclassical Néel states, which agrees with the exact results.
@@ 3.2 Antiferromagnetic Heisenberg Model on Triangular Lattices
We apply the DMMFT to another illustrative example, the AFHTL. The Hamiltonian for the AFHTL involving nearest neighbor anisotropic exchange interactions is given by
$\begin{split}H_{AFH}=&\sum_{i,j}\left[J_{xy}\left(S_{i}^{x}S_{j}^{x}+S_{i}^{y}S_{j}^{y}\right)+J_{z}S_{i}^{z}S_{j}^{z}\right]\\ =&\sum_{i,j}J\left[\mathcal{A}\left(S_{i}^{x}S_{j}^{x}+S_{i}^{y}S_{j}^{y}\right)+S_{i}^{z}S_{j}^{z}\right],\end{split}$ (21)
where $J_{z}$ and $J_{xy}$ are strengths of longitudinal and transverse exchange interactions. For antiferromagnetic couplings, it is more convenient to parameterize the model with $J=J_{z}>0$ and $\mathcal{A}=J_{xy}/J_{z}$. The phases of the system only depend on the dimensionless parameter $\mathcal{A}$ that specifies the exchange anisotropy. Our focus is on the easy-axis case, i.e., $0\leq\mathcal{A}\leq 1$. The two limiting points of $\mathcal{A}$ are: 1) $\mathcal{A}=0$, where the transverse exchange terms vanish (often referred to as the Ising limit), and 2) $\mathcal{A}=1$, where the exchange interactions have full rotational symmetry (often referred to as the Heisenberg limit). Both of the limits limits have been extensively studied in the literatures. Particularly, in the Ising limit, quantum fluctuations are additionally introduced by the transverse magnetic field, and the sign problem of the diagonal (longitudinal) frustrations can be mitigated in the QMC simulations [47, 49]; while in the Heisenberg limit the system orders in the so-called $120^{\circ}$-Néel state, as confirmed by various numerical methods [31, 32, 33, 41, 42].
While the phase diagram is believed to be bounded by the Ising and Heisenberg limits for a generic anisotropy parameter $10<\mathcal{A}<$ 1, computing the details of quantum phases becomes challenging, particularly due to the QMC sign problem when $\mathcal{A}\neq 0$. A study in Ref. 42 utilized large-size cluster mean-field theory (belonging to the class of conventional MFTs) with a scaling scheme, which provides a reasonable benchmark for the quantum phase diagram. With DMMFT, we can more efficiently and quantitatively capture the effects of quantum fluctuations with a small cluster of three sublattice sites.
For illustrative purposes, we focus on the phases of the AFHTL with easy-axis anisotropy subjected to a longitudinal external field, i.e.,
@@
where $H_{Z}=-\sum_{i}g\mu_{B}h_{z}S_{i}^{z}$ is the Zeeman coupling of the spins to the magnetic field $h_{z}$. In our DMMFT calculations, we set the gyromagnetic factor $g\mu_{B}=1$ and $h_{z}$ is measured in the corresponding natural unit.
Four phases emerge in the AFHTL with generic easy-axis anisotropy under a longitudinal external field. They are the coplanar ’“Y”-shaped phase, the collinear up-up-down (UUD) phase, the coplanar “V”-shaped phase, and the collinear polarized phase in the order of increasing $h_{z}$. All these phases are compatible with a cluster partition of three-sublattice sites and translational symmetries. A simple order parameter distinguishing these phases is the magnetization along the easy axis, i.e.,
$\begin{split}M^{z}_{c}=\sum_{\alpha\in c}S_{\alpha}^{z}=S_{A}^{z}+S_{B}^{z}+S_{C}^{z},\end{split}$ (23)
where the subscript $\alpha=A,B,C$ labels three sublattice sites within the cluster $c$. The polarized and the UUD phases are characterized by magnetization plateaux at $M^{z}_{c}=M_{z}^{c,(s)}$ and $M^{z}_{c}=M_{z}^{c,(s)}/3$, respectively, where $M^{z}_{c,(s)}=\frac{3}{2}$ is the saturation magnetization. The “V”-shaped phase is sandwiched in between the UUD and the polarized phases, while the “Y”-shaped phase appears at low fields. A distinguishing order parameter characterizing the non-collinearity of the “Y”-shaped phase is the vector chirality
@@
In our DMMFT calculations, we choose an extended cluster $\bar{c}$ comprising of four three-sublattice clusters with periodic boundary conditions. Utilizing periodic boundary conditions is to minimize the boundary effects, since the Hilbert subspaces selected by the reduced DM is sensitive to the boundary conditions as known in DMRG [50], While all four phases are homologous to their corresponding semiclassical ordered states, the quantum fluctuations play crucial roles in the “Y”-shaped and UUD phases. Specifically, quantum fluctuations lift the accidental degeneracies in the semiclassical ground state manifold, which is known as quantum order-by-disorder [45, 46], and significantly renormalize the magnitudes of the ordered spins [31, 32, 33].
@@
In Fig. 2 we present the dependence of the vector chirality, the magnetization, and the entanglement entropy of the ground states of AFHTL (with anisotropy parameter $\mathcal{A}=0.9$) on the longitudinal magnetic field, calculated with the DMMFT (red) and the conventional MFT (blue). The cut-off parameter $n_{c}$, as discussed in Sec. 2.3 interpolating the conventional MFT and the DMMFT, is set to $n_{c}=4$ for the DMMFT and $n_{c}=1$ for the conventional MFT.
@@
In Fig. 2(a), the vector chirality calculated with conventional MFT decreases monotonically as $h_{z}$ increases in the “Y”-shaped phase and vanishes in the UUD and the “V”-shaped phases. In sharp contrast, the vector chirality calculated with DMMFT, does not decrease monotonically in the “Y”-shaped phase. Notably, the vector chirality at $h_{z}=0$ in DMMFT is only about $30\%$ of that in conventional MFT. The discrepancy reflects the renormalization of the ordered spins due to the quantum fluctuations. Examining the magnitude of the ordered spins more closely , we find $|\langle\mathbf{S}\rangle|=0.269\pm 0.014$ for small anisotropy $\mathcal{A}=0.9$, This value is only about $50\%$ of $S=\frac{1}{2}$ but agrees with the expectation from previous reported values [31, 32, 33] . Since $\bm{\kappa}_{V}$ scales as $S^{2}$, a reduction factor of about $25\%$ is expected for the magnitude of the vector chirality at $h_{z}=0$. The curves of the DMMFT and the conventional MFT converge as $h_{z}$ increases and both vanish beyond $h_{c1}$ (the critical field separating the “Y”-shaped and the UUD phases determined according to the magnetization curves). With stronger $h_{z}$, the spins are aligned closer to the easy axis, and the effect of the quantum fluctuations weakens, which is clearly evidenced from the monotonically decreasing entanglement entropy in the “Y”-shaped phase as shown in Fig. 2(c). Therefore, the suppression of the quantum fluctuations and the aligning of the spins towards the collinear UUD configuration by the longitudinal magnetic field jointly lead to the non-monotonic dependence of the vector chirality $|\bm{\kappa}_{V}|$ on $h_{z}$.
As $h_{z}$ surpasses $h_{c2}$, the system transitions towards the “V”-shaped phase, deviating from the $\frac{1}{3}$ magnetization plateau of the UUD phase. Fig. 2(b) demonstrates a discernible difference in the critical field determined from the magnetization curves in DMMFT ($h^{\prime}_{c2}$) and in convetional MFT ($h_{c2}$). Notably, the range of the field stabilizing the UUD phase in DMMFT is larger than that in conventional MFT. Although the UUD configuration is homologous to the semiclassical correspondence,quantum fluctuations aid in stabilizing the UUD phase [42]. Especially in the Heisenberg limit, the UUD phase is only stable at one point of $h_{z}=h_{c2}=\frac{1}{3}h_{z}^{(s)}=1.5$ (where $h_{z}^{(s)}$ is the saturation field) in the classical phase diagram, while it extends to a finite range of the field in the quantum phase diagram. A similar effect of quantum fluctuations is also expected when $\mathcal{A}<1$. Since $n_{c}$ used in DMMFT is larger than that in conventional MFT, more quantum fluctuations are included, a larger field range ($h^{\prime}_{c2}>h_{c2}$) that stabilizes the UUD phase is expected.
@@
Additionally, a self-consistency check for the choice of $n_{c}=4$ in DMMFT can be confirmed from the entanglement entropy in the “Y”shaped phase. At $h_{z}=0$, we have $SE_{c}\approx 1.8\ln(2)$. Therefore, $\tilde{D}^{E}_{c}=\lceil\exp({SE}_{c})\rceil=4$.

### v2

@@ 3.1 Affleck-Kennedy-Lieb-Tasaki Model
The AKLT model is an exactly solvable system, showcasing non-trivial topological order. Notably, it features spin-$1/2$ (fractional to spin-$1$) edge states and exhibits a $4$-fold degeneracy of the ground states for an open chain [43], which persists even for $\beta<\frac{1}{3}$ deviating from its integrable point, provided the gap does not close [68, 60]. Moreover, the Néel order vanishes in the ground states. Therefore, the conventional MFTs fail to distinguish it from a trivial paramagnetic phase. Nevertheless, with the help of reduced DM and entanglement spectra, DMMFT is able to probe the topological order borne in its ground state.
@@
We apply the DMMFT to the AKLT model, focusing on the two-site cluster $c$ and solving the reduced DM $\rho_{c}^{(2)}$ self-consistently according to the mean-field equations as described in Sec. 2.2. Specifically, the reduced DM is calculated over an extended cluster encompassing the left and right nearest-neighboring two-site clusters to $c$. The associated Hilbert space is denoted as ~$\tilde{\mathscr{H}}_{\bar{c}}=\tilde{\mathscr{H}}_{l}\otimes\mathscr{H}_{c}\otimes\tilde{\mathscr{H}}_{r}$, where ~$\tilde{\mathscr{H}}_{l,r}\sim\Pi_{c}(\mathscr{H}_{c})$ are the Hilbert subspaces selected by the reduced DM, as per Eq.(13). We choose the cut-off parameter $n_{c}=4$.
In Fig. 1, the spectra of the two-site and single-site reduced DMs are presented. Blue circles represent the exact values, while red dots depict the DMMFT results. The spectrum of the two-site reduced DM obtained with the DMMFT reasonably aligns with the exact values [Fig. 1(a)], notably capturing the correct degeneracies. The entanglement entropy of the two-site cluster evaluated by DMMFT is $SE_{c,\text{MF}}^{(2)}=1.58$. Although it is slightly larger than the exact value $SE_{c,\text{exact}}^{(2)}=1.3689$, this discrepancy is expected due to the cut-off parameter $n_{c}$ chosen being marginally larger than $\exp(SE_{c,\text{exact}}^{(2)})=3.9310$. Fig. 1(b) displays the spectrum of the single-site reduced DM, where the DMMFT results align excellently with the exact values. The single-site entanglement entropy $SE_{c}^{(1)}=1.0986=\ln(3)$ also matches. Furthermore, the expectation values of two spins within the cluster $\langle S_{1,2}^{\alpha}\rangle=\text{Tr}_{c}(S_{1,2}^{\alpha}\rho_{c})$ both vanish. This implies that the ground states found by DMMFT are not biased towards the semiclassical Néel states, which agrees with the exact results.
@@ 3.2 Antiferromagnetic Heisenberg Model on Triangular Lattices
We apply the DMMFT to another illustrative example, the simple AFHTL. The Hamiltonian for the AFHTL involving nearest-neighbor anisotropic exchange interactions is given by
$\begin{split}H_{AFH}=&\sum_{(i,j)}\left[J_{xy}\left(S_{i}^{x}S_{j}^{x}+S_{i}^{y}S_{j}^{y}\right)+J_{z}S_{i}^{z}S_{j}^{z}\right]\\ =&\sum_{(i,j)}J\left[\mathcal{A}\left(S_{i}^{x}S_{j}^{x}+S_{i}^{y}S_{j}^{y}\right)+S_{i}^{z}S_{j}^{z}\right],\end{split}$ (21)
where $J_{z}$ and $J_{xy}$ are strengths of longitudinal and transverse exchange interactions. For antiferromagnetic couplings, it is more convenient to parameterize the model with $J=J_{z}>0$ and $\mathcal{A}=J_{xy}/J_{z}$. The phases of the system only depend on the dimensionless parameter $\mathcal{A}$ that specifies the exchange anisotropy. Our focus is on the easy-axis case, i.e., $0\leq\mathcal{A}\leq 1$. The two limiting points of $\mathcal{A}$ are: 1) $\mathcal{A}=0$, where the transverse exchange terms vanish (often referred to as the Ising limit), and 2) $\mathcal{A}=1$, where the exchange interactions have full rotational symmetry (often referred to as the Heisenberg limit). Both limits have been extensively studied in the literatures. Particularly, in the Ising limit, quantum fluctuations are additionally introduced by the transverse magnetic field, and the sign problem of the diagonal (longitudinal) frustrations can be mitigated in the QMC simulations [58, 69]. Meanwhile, in the Heisenberg limit, the system orders in the so-called $120^{\circ}$-Néel state, as confirmed by various numerical methods [44, 45, 46].
While the phase diagram of the AFHTL in a longitudinal external magnetic field is believed to be bounded by the Ising and Heisenberg limits for a generic anisotropy parameter $10<\mathcal{A}<$ 1, computing the details of quantum phases remains challenging, particularly due to the QMC sign problem when $\mathcal{A}\neq 0$. In previous studies such as those in Refs. 52, 53, the magnetization has been examined using large-size cluster mean-field theory (belonging to the class of conventional MFTs) with a scaling scheme. Additionally, Ref. 47 and Ref. 49 have reported studies of the magnetization with ED, respectively, and Ref. 48 has presented phase diagram obtained with DMRG  [^1] We thank the anonymous referee for pointing Refs. 47, 49, 48, 50 to us. Cf. Ref. 50 and references therein for earlier studies.. These previous studies provide comprehensive understandings of the AFHTL in a magnetic field and serve as reasonable benchmarks for the quantum phase diagram.
Below, we present our investigation of the AFHTL using DMMFT. For benchmarking purposes, we also performed a calculation with the state-of-the-art technique DMRG. We compare our results obtained with DMMFT to those from DMRG as well as results reported in the literatures. Remarkably, we find quantitative agreement between the results obtained with DMMFT and those from the more sophisticated methods. This alignment is achieved with a small cluster of just three lattice sites in DMMFT, demonstrating the efficiency of the method.
The Hamiltonian describing the AFHTL in a longitudinal external field includes both the exchange interactions and the Zeeman couplings,
@@
where $H_{Z}=-\sum_{i}g\mu_{B}h_{z}S_{i}^{z}$. In our DMMFT calculations, we set the gyromagnetic factor $g\mu_{B}=1$ and $h_{z}$ is measured in the corresponding natural unit.
Four phases emerge in the AFHTL with generic easy-axis anisotropy under a longitudinal external magnetic field. They are the coplanar ’“Y”-shaped phase, the collinear up-up-down (UUD) phase, the coplanar “V”-shaped phase, and the collinear polarized phase, in the order of increasing $h_{z}$. All these phases are compatible with a cluster partition of three lattice sites and translational symmetries. A simple order parameter distinguishing these phases is the magnetization along the easy axis, i.e.,
$\begin{split}M^{z}_{c}=\sum_{i\in c}S_{i}^{z}=S_{A}^{z}+S_{B}^{z}+S_{C}^{z},\end{split}$ (23)
where the subscript $i=A,B,C$ labels three sublattice sites within the cluster $c$. The polarized and the UUD phases are characterized by magnetization plateaux at $M^{z}_{c}=M^{z}_{c,(s)}$ and $M^{z}_{c}=M^{z}_{c,(s)}/3$, respectively, where $M^{z}_{c,(s)}=\frac{3}{2}$ is the saturation magnetization of the cluster. The “V”-shaped phase is sandwiched in between the UUD and the polarized phases, while the “Y”-shaped phase appears at low fields. A distinguishing order parameter characterizing the non-collinearity of the “Y”-shaped phase is the vector chirality
@@
Despite that all four phases are homologous to their corresponding semiclassical Néel ordered states, the quantum fluctuations play crucial roles in the “Y”-shaped and UUD phases. Specifically, quantum fluctuations lift the accidental degeneracies in the semiclassical ground state manifold, which is known as quantum order-by-disorder [56, 57], and significantly renormalize the magnitudes of the ordered spins [44, 45, 46] as well as the order parameters.
@@
In our DMMFT calculations, we choose an extended cluster $\bar{c}$ comprising of four three-lattice clusters with periodic boundary conditions. Employing the periodic boundary conditions is to minimize the boundary effects, since the Hilbert subspaces selected by the reduced DM is sensitive to the boundary conditions as known in DMRG [70]. The geometry and implementation details of DMMFT for the AFHTL are elaborated in Append. B, along with provided pseudocodes.
@@
In Fig. 2, we present the dependence of the vector chirality, the magnetization, and the entanglement entropy of the ground states of AFHTL (with anisotropy parameter $\mathcal{A}=0.9$) on the longitudinal magnetic field, calculated with the DMMFT (red) and the conventional MFT (blue). The cut-off parameter $n_{c}$, as discussed in Sec. 2.3 interpolating the conventional MFT and the DMMFT, is set to $n_{c}=4$ for the DMMFT and $n_{c}=1$ for the conventional MFT.
@@
In Fig. 2(a), the vector chirality calculated with conventional MFT decreases monotonically as $h_{z}$ increases in the “Y”-shaped phase and vanishes in the UUD and the “V”-shaped phases. In sharp contrast, the vector chirality calculated with DMMFT, does not decrease monotonically in the “Y”-shaped phase. Notably, the vector chirality at $h_{z}=0$ in DMMFT is only about $30\%$ of that in conventional MFT. The discrepancy reflects the renormalization of the ordered spins due to the quantum fluctuations. Upon closer examination of the magnitude of the ordered spins, for a small anisotropy $\mathcal{A}=0.9$, we find $|\langle\mathbf{S}\rangle|=0.269\pm 0.014$. This value, only about $50\%$ of $S=\frac{1}{2}$, aligns with expectations based on previously reported values [44, 45, 46]. Given that $\bm{\kappa}_{V}$ scales as $S^{2}$, a reduction factor of approximately $25\%$ is expected for the magnitude of the vector chirality at $h_{z}=0$. The curves of the DMMFT and the conventional MFT converge as $h_{z}$ increases, both vanishing beyond $h_{c1}$ (the critical field separating the “Y”-shaped and the UUD phases determined according to the magnetization curves). With stronger $h_{z}$, the spins are aligned closer to the easy axis, and the effect of the quantum fluctuations weakens, which is clearly evidenced from the monotonically decreasing entanglement entropy in the “Y”-shaped phase as shown in Fig. 2(c). Therefore, the suppression of the quantum fluctuations and the alignment of the spins towards the collinear UUD configuration by the longitudinal magnetic field jointly lead to the non-monotonic dependence of the vector chirality $|\bm{\kappa}_{V}|$ on $h_{z}$.
[…]
@@
Fig. 3(b) shows the magnetization curve calculated with DMRG for YC6666-$30$ and bond dimension $D_{b}=400$ (black circles), overlaid with the results of conventional MFT (blue dots) and DMMFT (red dots). The DMRG results are in excellent agreement with the DMMFT results, confirming the nonlinear magnetization in low field. Additionally, a self-consistency check for the choice of $n_{c}=4$ in DMMFT can be confirmed from the entanglement entropy in the “Y”-shaped phase. At $h_{z}=0$, we have $SE_{c}\approx 1.8\ln(2)$. Therefore, $\tilde{D}^{E}_{c}=\lceil\exp({SE}_{c})\rceil=4$.
On the other hand, as $h_{z}$ surpasses $h_{c2}$, the system transitions towards the “V”-shaped phase, deviating from the $\frac{1}{3}$ magnetization plateau of the UUD phase. Fig. 2(b) demonstrates a discernible difference in the critical field determined from the magnetization curves in DMMFT ($h^{\prime}_{c2}\approx 2.2$) and in conventional MFT ($h_{c2}\approx 2.0$). Notably, the range of the field stabilizing the UUD phase in DMMFT is larger than that in conventional MFT. Although the UUD configuration is homologous to its semiclassical correspondence, quantum fluctuations contribute significantly to stabilize the UUD phase [52]. Especially in the Heisenberg limit, the UUD phase is stable only $h_{z}=h_{c2}=\frac{1}{3}h_{z}^{(s)}=1.5$ ($h_{z}^{(s)}$ is the saturation field) in the classical phase diagram, whereas it extends to a finite field range due to quantum fluctuations in the quantum phase diagram. A similar effect of quantum fluctuations is also anticipated when $\mathcal{A}<1$. Since $n_{c}$ used in DMMFT is larger than that in conventional MFT, which includes more quantum fluctuations, a larger field range ($h^{\prime}_{c2}>h_{c2}$) that stabilizes the UUD phase is expected.
To provide a quantitative comparison, we further compare the value of $h_{c2}$ obtained from DMMFT to those estimated from our DMRG results and those inferred from the literatures. The inset of Fig. 3(b) provides a zoomed-in view of the transition between the UUD phase and the “V”-shaped phase. Our DMRG results indicate $h_{c2}^{*}\approx 2.06$, which falls between the $h_{c2}$ of conventional MFT and $h^{\prime}_{c2}$ of DMMFT. Few precise values of $h_{c2}$ for $\mathcal{A}=0.9$ have been reported in the literatures [47, 48, 49, 52, 53]. Inferring from the reported values of $h_{c2}$ for the isotropic Heisenberg model, we conclude that the $h_{c2}^{*}$ estimated from our DMRG results may have slightly underestimated the critical field, probably due to the finite-size effect of $N_{y}=6$ along the transverse direction of the cylinder [compared to Fig. 1 in Ref. 48 and Fig. 7(a) in Ref. 53 for scaling to thermodynamic limit]. Whereas $h_{c2}^{\prime}$ estimated from our DMMFT results may have slightly overestimated the critical field due to our choice of $n_{c}=4$, which is larger than the speculated optimal $n_{c}^{*}=2$ estimated from the entanglement entropy. This leads to an inclusion of too many quantum fluctuations favoring the UUD phase. Moreover, as $h_{z}$ drives the system deeper inside the “V”-shaped phase, the magnetization curve obtained from our DMRG calculation gradually converges towards that of DMMFT.
In conclusion, our study of the AFHTL highlights the systematic improvement made by DMMFT over the conventional MFT. Remarkably, even with just a small cluster of three lattice sites, DMMFT achieves precision comparable to that of large system sizes calculated using DMRG or conventional MFT plus scaling. This underscores the efficiency and efficacy of DMMFT in capturing the effects of quantum fluctuations.